# Decentralization of Ethereum’s Builder Market

### v3

@@ 5.1 Pivotal Private Order Flow Providers
For each pivotal provider, we compute the fraction of auctions in which it is pivotal. We call this metric the pivotal level, a number from 0 to 1. Fig. 8 plots the pivotal level of the top-5 providers. A builder who lacks access to any of their order flows will lose the majority of auctions. However, we found that order flows from these providers are inaccessible to all builders, which requires further investigation. Other observations from Fig. 8 are provided in Appx. F.2.
5.2 Trust Barrier
A direct reason for the access barrier is that private order flow providers are at a disadvantage when interacting with builders and must trust builders not to unbundle the transactions. However, in practice, builders occasionally unbundle searcher’s bundles. For example, the f1b builder unbundled a bundle from the searcher jaredfromsubway.eth at block height 18476515. Although the searcher was compensated afterward, this incident highlights the fragile security foundation of the MEV-Boost auction.
Due to the lack of trustworthy fair exchange mechanisms, providers have to rely on an informal reputation system for self-protection, sending order flows only to builders who meet their requirements. To understand the access barrier caused by the trust crisis, we investigate different providers and their requirements for accessing order flows.
Searchers.   Since searchers are operated by anonymous entities without any public communication channels, there is no public statement about which builders may access their flows. Existing studies indicate that searchers tend to send their order flow to the top builders [81] or share it exclusively with vertically integrated builders [81, 80, 40].
Telegram trading bots.   The requirements for Telegram trading bots, in terms of sending order flow to builders, depend on whether they utilize channels like MEV Blocker [56, 48]. If so, their requirements are aligned with those of these channels. Banana Gun does not explicitly state the requirements for builders but we observe its preference for Titan, which we will discuss later.
Channels.   Since channels are public-facing services, we were able to extract their access requirements from the documentation and/or through their communication channels. The requirements of each provider are summarized in Tab. 5.
The channels impose various requirements on a builder’s reputation. Market share—the percentage of blocks built by a given builder—is a metric most relay providers use to evaluate builders’ reputations. Intuitively, a builder with a high market share, seen as reputable, is less likely to compromise private order flows as doing so will forgo their future revenue. MeowRPC [58] sends the order flow to trusted builders with high market share. Similarly, MEV-Share [37] requires that a builder possesses a competitive market share. MEV Blocker [60] requires that a builder must maintain at least 1% market share over each week of one month.
Subsidization.   New builders face a “chicken and egg” problem as they need access to private order flows to win auctions and gain market share, but the private order flow providers only serve builders with a decent market share. In practice, builders pay out of pocket to overbid in MEV-Boost auctions, a practice known as block subsidization [36], representing an entry barrier for new builders. We quantify the minimal cost for a new builder to meet the MEV Blocker’s requirements (Appx. F.5), which reflects the costs of building a reputation. The minimal cost continues to increase and can be up to 7.4 ETH. This cost represents a lower bound since we do not have the corresponding numbers for other providers. Note that although subsidization can help meet the requirements set by channels, it cannot ensure that searchers’ trust concerns are addressed, allowing builders to receive their order flows.
5.3 Incentive for Integration
Another access barrier to private order flows is that providers exclusively share order flows with specific builders, known as integration [40]. We compute each provider’s builder-specific pivotal level to understand the relationship between providers and builders. For a given provider $P$, its pivotal level specific to builder $B$ is the pivotal level of the MEV-Boost auctions where $B$ is the winner.
Figure 9: Weekly builder-specific pivotal level for top-3 builders. MEV-Blocker and MEV-Share are hidden for clarity.
Fig. 9 plots selected providers’ builder-specific pivotal levels for the top three builders, ranked based on their weekly market share in July 2024. We include providers whose weekly pivotal levels exceed 25% for any of the three builders over two months to illustrate these builder-specific pivotal levels. MEV Blocker and MEV-Share are excluded from this analysis, as all builders connected to them reportedly receive the same order flows. We note that some providers have a pivotal level exceeding 50% only with specific builders, suggesting a possible collaborative relationship. The searcher 0xA69b… consistently influences over 50% of the MEV-Boost auctions won by Beaver, while the searchers 0x0087…, 0x2800…, and 0x51c7… successively affect those won by Rsync. These pivotal providers are all integrated searchers of the corresponding builders that have been identified in previous works [42, 40]. This observation corroborates findings from previous studies [42, 40, 80], confirming that the order flows from integrated searchers are crucial to those builders.
Furthermore, we observe a clear variation in the importance of providers for different builders. For example, the significance of Banana Gun to Titan is greater than it is to other builders. We further investigate the relationship between Banana Gun and Titan and suppose that Banana Gun only sends valuable order flow to Titan when MEV is significant. A case in this point is slot 8614161, where only Titan received the valuable order flow from Banana Gun. The much lower bids from other builders suggest they didn’t receive this order flow; otherwise, they could have outbid Titan. We illustrate this new integration pattern in Appx. F.3. The results in Fig. 15 indicate that this integration pattern also applies to jaredfromsubway.eth and Beaver. Both integrations emerged around January 2024.
There is a clear incentive for integration between providers and builders. If MEV from a provider can determine the auction outcome, sending order flow to an integrated builder is better than sending it to multiple builders. The integrated builder only needs to bid slightly higher to win, and they can share the surplus MEV. However, this integration is detrimental to proposers, as it exacerbates the inequality in builders’ block-building capacities and can eventually lead to significant losses for them.
Finding IV: Two providers (Banana Gun and jaredfromsubway.eth) exclusively share high-value order flows with specific builders (Titan and Beaver) respectively, suggesting a new integration pattern.
@@ 6 Related Work
Empirical study on Ethereum’s builder market.   Several previous works studied the centralization of the builder market shares. An early exploration by Yang et al. [92] analyzed the market share of relays and builders. Their findings indicated a centralization of Ethereum’s builder market, with Flashbots and builder0x69 accounting for over 53% of the market share between September 15th, 2022, and November 30th, 2022. Similar studies on the builder market conducted by [84, 41] also confirmed the centralization in the builder market. Their observations were based on the market share and the distribution of total profits generated within the PBS ecosystem. Moreover, several online dashboards [87, 49, 7, 36] offer a dynamic view of the ecosystem’s trends by presenting real-time PBS analytics. However, these studies do not quantify the causes of centralization or the security implications of a centralized builder market.
@@
Private order flow.   Many studies also discuss the importance of private order flows, particularly focusing on private order flows from integrated searchers. Thiery [80] counted the total number of transactions, along with their values, from June 1 to July 15, 2023. He found that private order flows account for only 30% of the transactions yet represent 80% of the total value paid to builders. Our study highlights a similar result: in over 50% of blocks, private order flows account for more than 60% of the block value. Gupta et al. [40] confirmed that builders are more likely to win the MEV auction when their integrated searchers can provide high-value exclusive order flow by effectively exploiting arbitrage between a centralized exchange (CEX) and a decentralized exchange (DEX). The dashboard by Winnie [91] and the study by Titan [81] analyzed the vertical integration between searchers and builders through the quantity and value of transactions sent from searchers to builders. Heimbach et al. [42] observed that during their study, Beaver Builder received a total of 1,941.1 ETH in transaction fees but paid proposers 6,620.94 ETH. This large gap suggests that Beaver might have received significant off-chain profits from its integrated searchers. These studies show that three integrated builders (Beaver, Rsync, and Manta) gain a significant advantage in order flow from their integrated searchers. Our study made the same observation and further quantified the pivotal level.
@@
Strategic behaviors in MEV-Boost auctions.   The strategies used by different entities in the MEV-Boost auctions and their potential impacts have attracted significant interest from researchers. The strategic behaviors of proposers were first studied by [76, 84] focusing on the timing game where proposers strategically delay block proposals to maximize their profits. Öz et al. [66] further conducted an agent-based simulation to analyze how waiting games affect consensus stability, finding that a delay strategy can be profitable and not degrade consensus if sufficiently many validators adopt it. A recent study [1] empirically investigated timing games in the wild and how block proposal delays affect the block fork rate. The results indicate that three validator sets–P2P, Kiln, and Attestant, that are known or suspected of playing timing games–did not delay their block proposals enough to cause their blocks to be forked. Wu et al. [90] proposed a game-theoretic model for MEV-Boost auctions, using simulations to examine various builders’ bidding strategies, including naive, adaptive, last-minute, and bluff bidding. Their results highlighted the importance of latency and exclusive order flow in the effectiveness of bidding strategies. Pai et al. [67] analyzed the latency advantage of searcher-builder integration in adjusting bidding values. The strategies discussed in these papers highlight the potential factors that can make MEV-Boost auctions uncompetitive or inefficient, while our paper quantifies these uncompetitive and inefficient auctions.
@@
Causes and implications of centralized builder market.   A provider may exclusively share its order flow with a single builder. Previous works such as [46, 38, 40] pointed out that exclusive order flows have potential centralizing effects on the builder market, because builders with exclusive order flows are more likely to win MEV-Boost auctions, thus gaining further advantages in receiving order flows from non-integrated searchers and in winning order flow auctions. The decentralization of Ethereum’s builder market not only benefits itself but also positively impacts the decentralization of validators. A recent study showed that heterogeneity among validators in building blocks may lead to centralization; however, a decentralized and competitive builder market can encourage validator decentralization [4]. We complement these theoretical analyses with empirical studies to analyze potential problems caused by the current centralized builder market.
Blockchain decentralization metrics.   Nakamoto coefficient was proposed to describe the minimum number of entities required to disrupt the blockchain’s subsystem [77]. Other metrics, such as the Gini coefficient [18], Shannon entropy [51], and Herfindahl–Hirschman index [73], have also been utilized to measure decentralization (or the lack of it) in staking distribution [39, 47], mining power [52, 89, 47], and block building [41]. An SoK paper [93] comprehensively summarized the existing works on blockchain centralization; however, none explored decentralization from the perspective of entry barriers or auctions.
7 Discussion
7.1 Mitigation
Our results suggest that proposers suffer from significant losses because of builder centralization, which can lead to two undesirable outcomes: First, PBS could become unstable. Second, MEV measured by the auctions may be inaccurate.
In this section, we propose a structural change to the existing MEV supply chain Sec. 2.2 and a solution based on the new supply chain structure to address these two problems.
In the new supply chain, only one builder receives order flows from providers, and proposers can only accept blocks from this builder.
Protocol.   Our solution involves having the builder run the truthful mechanism in [57] in a Trusted Execution Environment (TEE) [15] The pseudocode for the TEE builderis described in Fig. 17.
The builder provides two functions. First, it opens RPC endpoints to receive the private order flows from the providers. Providers send their order flows along with a bid indicating the MEV in their order flows. Second, when it receives a signal for block building (e.g., a new slot starts in Ethereum), it will run an ordering algorithm to build a block with optimal MEV based on the received order flows (a reasonably performant ordering algorithm is described in Appx. E).
The TEE builder runs the truthful mechanism in [57]: it refunds all providers for their marginal value contribution to the block. For each included order flow, the builder builds an alternative block without that specific order flow to calculate its marginal value contribution by comparing the MEV extracted between these two blocks. Next, we outline the key advantages of this solution.
Security guarantees.   TEEs offer a secure and isolated environment, ensuring that private order flows remain confidential and protected from external tampering. Moreover, order flow providers can utilize remote attestation to verify that the builder running within the TEE is indeed the intended one, thereby effectively mitigating trust concerns.
Incentive compatibility.   [57] shows that for providers, truthfully bidding is a dominant strategy when they are refunded based on their marginal value contribution. The formal proof and the required assumptions are in [57]. Here, we offer an intuitive explanation: a provider effectively pays the second price for their order flows when refunding their marginal value contribution.
Analysis.   Since truthfully bidding is an incentive-compatible dominant strategy for all providers, the TEE builder can compute the total MEV in a block from the providers’ bids, achieving an accurate MEV oracle. Moreover, because the builder program in the TEE can be programmed to pay the proposer fairly, the instability of PBS is also addressed.
Future works.   We leave the analysis of whether the new supply chain structure is sustainable in the long term for future work. Intuitively, since providers are incentivized to send order flows only to the builder due to refunds, other builders will naturally be driven out of the market, leaving proposers to only accept blocks from the TEE builder. Furthermore, since real-world proposers might behave rationally, exploring the new equilibrium that emerges among a rational provider, builder, and proposer is another intriguing question to be addressed. Finally, this solution provides an accurate MEV oracle and transparent MEV information. These insights open up new possibilities for future MEV mitigation directions.
7.2 Limitations of Our Approach and Data
True value.   We compute the true values of bids using a similar approach as that in MEV analysis tools [20, 31] and previous works [41, 80]. This approach has a few limitations. First, it relies on the knowledge of the set of addresses controlled by each builder. Most builders use well-known addresses to receive payments from searchers and make payments to proposers, but it is possible that some builders use less publicized addresses. We cross-validated our builder address dataset with other sources to ensure completeness (see Sec. 3.1), but if we missed a builder’s address, the computation of extracted value (and hence true value) would be inaccurate. Secondly, this approach neglects off-chain profits, such as off-chain payments received from searchers. Off-chain payments are more likely to happen between integrated builder-searcher pairs. We might underestimate the true value of a block due to unknown off-chain payments, which could lead to the misclassification of an efficient auction as inefficient due to subsidization. However, in our study, we did not observe any excessively large subsidies. Furthermore, underestimating the winner’s true value does not affect our assessment of competitiveness. We might misclassify an uncompetitive auction as competitive if we underestimate the second-highest true value. However, such underestimation should be uncommon, as we observe that most builders bid around our computed true value. It would be irrational for builders not to bid higher to win the auction if they actually have a higher true value.
Missing pivotal providers.   Our results in Sec. 5.1 might omit some pivotal providers since the providers we can identify are a subset of all providers. For example, BloXroute (a widely used channel) does not publish the list of transactions they processed. Thus, we can’t evaluate its pivotal level. This limitation is common in existing studies [53, 54]. However, there will not be false positives (classification of a non-pivotal provider as pivotal) because we identify a given transaction’s provider based either on the official dataset from MEV-Share and MEV Blocker or on the specific searchers’ addresses (see Sec. 3). Thus, the significance of each provider will only be underestimated rather than overestimated.
7.3 Security Problems in PBS
As we mentioned in Sec. 5.2, one reason for the access barrier to private order flows is the difficulty of establishing trust. A malicious builder, relay, or proposer may attack providers. A real-world incident involved a malicious proposer who double-signed a header to unbundle a builder block and further stole MEV from searchers by proposing a different block [61].
Another security concern is censorship. Ethereum’s censorship resistance currently depends heavily on whether builders choose to censor, particularly the actions of the top builders. For example, before the top builder, Beaver, stopped censoring in July 2024, over 50% of blocks were censoring [86]. Our study reveals that builder centralization remains difficult to overcome, potentially exacerbating the censorship problem in Ethereum.

### v4

@@ 5.1 Pivotal Private Order Flow Providers
Pivotal level.   For each pivotal provider, we compute the fraction of auctions in which it is pivotal. We call this metric the pivotal level, a number from 0 to 1. A higher pivotal level indicates a provider’s greater significance in the MEV-Boost auctions. Fig. 9 plots the pivotal level of the top-5 providers. A builder who lacks access to any of their order flows would lose the majority of auctions. However, order flows from the pivotal providers may not be accessible to all builders, which requires our further investigation. We also make three observations from Fig. 9, which are presented in Appx. E.1 due to space constraints.
Figure 9: The pivotal level of top-5 pivotal providers over time. A provider $P$ is pivotal if the winner of an auction could have lost without private order flows from $P$. A line may not start from the beginning because the provider did not exist at that point.
Accessibility by multiple builders.   Among the top five pivotal providers, MEV-Share and MEV Blocker are public-facing services with clear documentation and open communication channels. This enables us to analyze their statements regarding how their order flows are distributed to different builders—both claim that every connected builder receives the order flows [35, 36]. Although their order flows are not accessible to all builders (as summarized in Tab. V), top builders such as Beaver, Titan, and Rsync have access to them. Ideally, competition among these top builders will ensure that the value from these order flows is ultimately captured by the proposers.
Notably, while these order flows are distributed equally to all connected builders, inequality still exists. Order flow providers may exclude less reputable builders due to trust concerns. Although this exclusion may not harm proposers, it could exacerbate builder centralization, as new builders must expend significant effort to meet the requirements imposed by these pivotal providers. Due to page constraints, we provide a more detailed discussion of the trust barrier and how it prevents builder decentralization in Appx. E.2.
5.2 Exclusive Pivotal Providers
Pivotal providers like MEV-Share and MEV Blocker send their order flows to connected builders who satisfy their requirements. Many other providers (e.g., most searchers) do not publicly share their order flows; some only send order flows exclusively to specific builders. If multiple builders receive the same order flows, the proposer can, in principle, capture the value due to the competition. In contrast, a builder receiving exclusive order flows is detrimental to proposers, as it can exacerbate inequality and ultimately lead to losses for proposers.
The exclusive private order flows from a provider to a builder indicate a special relationship called integration [20]. In this section, we further study the integration between builders and providers in the current market and how much loss proposers incur due to the integration.
Case study of block 21,000,813.   To better understand exclusive order flows and their impact, we go through a case study of block 21,000,813, which has the highest estimated proposer loss in October 2024. In the auction for this block, Titan received a valuable order flow from Banana Gun (55.5 ETH) and bid 2.8 ETH, while the second-highest bid was only 0.028 ETH. The significantly lower bids from other builders suggest they did not receive this order flow; otherwise, they could have bid closer to Titan.
In this case, the order from Banana Gun dominated the auction, as whoever had that order could win. Thus, there is a clear incentive for the provider when its order flows are highly profitable: sending order flow to a select builder is more advantageous than distributing it to multiple builders. The select builder only needs to bid slightly higher to win, allowing the provider and the builder to share the surplus MEV (52.7 ETH in this case).
New integration pattern.   The above case highlights a potential integration between Titan and Banana Gun. Previous studies [20, 38] analyze an integration pattern between builders and searchers engaged in arbitrage between centralized exchanges (CEXs) and decentralized exchanges (DEXs). These searchers may be run by the same party as their integrated builders and exclusively send order flows to them, making the integrated builders more likely to win auctions when MEV from CEX-DEX arbitrage is significant.
Unlike known integrated searchers, Banana Gun is operated by a different party from Titan and typically sends order flows to multiple builders in most cases. For example, in October 2024, Banana Gun’s order flows were included in more blocks built by other builders than in those built by Titan. This suggests that Titan may only receive exclusive order flows from Banana Gun when those flows are particularly profitable. Similarly, other pivotal providers may tend to integrate with top builders when their order flows are highly profitable due to incentives.
To confirm this conjecture and evaluate the extent to which proposer loss is related to integration, we need to investigate exclusive order flows from providers to builders.
Challenges of identifying exclusive order flows.   Identifying exclusive order flows is challenging, as we do not know the exact order flows each builder receives. The on-chain blocks only record the winning bids of MEV-Boost auctions, which reflect the order flows received by the winners. Even though we have the full bids dataset from 2023, it remains unclear whether the builders did not receive an order flow or chose to exclude it. Analyzing the number and value of transactions sent from providers to builders is also unreliable because providers might selectively send exclusive order flows, and the difference may not be very significant.
Existing work [20] do not provide a methodology for identifying exclusive order flows, while [38] identifies integration by analyzing the correlation between the volume of searchers’ CEX-DEX arbitrage and the builders’ daily market share of blocks. This approach cannot be directly applied to identify generalized integration between providers and builders, as different providers may perform various MEV activities beyond CEX-DEX arbitrage.
Most importantly, since we are interested in how many proposer losses are related to integration, identifying a builder-provider integration pair alone cannot accurately answer this question. As discussed earlier, a provider may send order flows exclusively to one builder under specific conditions. Therefore, attributing all proposer losses in a builder’s blocks to its integration might be inaccurate.
before 2024 2024
Provider MEV Activity Builder Blocks (#) Losses (ETH) (%) Provider MEV Activity Builder Blocks (#) Losses (ETH) (%)
SCP CEX-DEX arbitrage Beaver 18,509 4,149.3  (21.0) Banana Gun Sniper Titan 33,748 19,401.3  (37.1)
Searcher: 0x57c1… CEX-DEX arbitrage Addr: 0x4737… 7,166 1,037.3  (5.2) SCP CEX-DEX arbitrage Beaver 10,087 3,249.8  (6.2)
Wintermute CEX-DEX arbitrage Rsync 6,555 553.1  (2.8) jaredfromsubway.eth Sandwich Beaver 20,213 2,575.4  (5.0)
Searcher: 0xf8b7… CEX-DEX arbitrage Manta 4,606 428.2  (2.2) Searcher: 0x6980… Atomic arbitrage I can haz block? 513 2,484.9  (4.8)
Searcher: 0x9dd8… Atomic arbitrage Addr: 0x3bee… 1,600 411.5  (2.6) Searcher: 0x6454… Liquidation Addr: 0x3bee… 793 1,229.6  (2.4)
Searcher: 0x761d… Atomic arbitrage builder0x69 12 287.0  (1.5) c0ffeebabe.eth Liquidation Ty For The Block 367 1,003.6  (1.9)
Searcher: 0x6454… Liquidation Addr: 0x3bee… 539 262.5  (1.5) Banana Gun Sniper Beaver 2,430 987.4  (1.9)
Searcher: 0x5e51… Atomic arbitrage eth-builder 266 223.2  (1.1) WINTERMUTE CEX-DEX arbitrage Rsync 6,503 976.8  (1.9)
Searcher: 0xeeaa… Atomic arbitrage eth-builder 51 194.1  (1.0) Maestro Sniper Beaver 10,103 520.3  (1.0)
Searcher: 0xe20c… Atomic arbitrage Addr: 0x5638… 579 168.5  (0.9) Searcher: 0x77d5… Atomic arbitrage Rsync 782 345.3  (0.7)
TABLE III: The top 10 providers sending exclusive order flows and their corresponding builders, contribute to the highest estimated proposer loss (before and in 2024).
Our approach.   Despite the challenges mentioned above, we are still able to infer whether an order flow is likely to be exclusively sent to a builder based on a reasonable assumption: all builders are rational and do not collude, so they will compete upon receiving the same order flow. We justify this assumption with our study in Sec. 4, which shows that over 90% of MEV-Boost auctions are competitive. Based on this assumption, we design an algorithm to infer if an order flow is exclusively sent to a builder.
For a given slot, we obtain the winning bid and compute the second-highest bid from our partial auction dataset. Then, for the winning bid in this slot, we identify the private order flows within this block and their associated providers using our private transaction dataset. For each provider, we evaluate whether the combined value of the private order flow from this provider and the public order flow exceeds the second-highest builder’s bid multiplied by a threshold. This threshold, set at 1.15, was derived from our analysis of historical MEV-Boost auctions, where less than 5% of builders withheld bids that exceeded this value. The threshold accounts for the possibility that builders may bid slightly below their true valuations. If this condition holds, it suggests that the provider exclusively sends order flows to the builder, thus giving the winner a competitive advantage.
Since there is no ground truth for exclusive order flows, we validate our method using four builder-searcher integrations identified in previous work [38]. Our algorithm cannot identify exclusive order flows when their value is insignificant, i.e., when they cannot determine the auction outcome. Thus, our approach will have false negatives. We focus on precision, specifically assessing, for each provider, the percentage of exclusive order flows correctly attributed to their well-known integrated builder.
For the known builder-searcher integrations—Beaver with SCP, Rsync with Wintermute, Manta with searcher 0xf8b7…, and anonymous builder 0x4737… with searcher 0x57c1…—our methodology achieves accuracies of 99%, 96%, 99%, and 99%, respectively. Except for Rsync and its integrated searcher, our methodology achieves very high precision: less than 1% of order flows are inaccurately identified as exclusive (false positive). Further investigation into the identified exclusive order flows of Wintermute shows that about 70% of the false positives occurred in blocks built before Rsync began block-building. At this point, Rsync might send order flows to multiple builders. However, only one builder included the order flow before the auction concluded. Besides the known integrations, we also identified previously unknown integrations involving the top-5 pivotal providers—Banana Gun, Maestro, and jaredfromsubway.eth—as listed in Tab. III.
Proposer loss due to integration.   For each builder-provider integration pair identified by our algorithm, we count the number of blocks where the builder exclusively receives the order flows and calculate the total estimated proposer loss in these blocks. Tab. III lists the top 10 providers and their corresponding builders receiving exclusive order flows, ranked by the resulting losses before and in 2024 respectively. We classify each provider’s MEV activity using previous studies [20, 38] and MEV analysis tools [59, 56]. For simplicity, we label each searcher with only its predominant MEV activity. Telegram bots, typical sniper bots [60], automatically assist users in quickly buying new tokens at launch by monitoring liquidity pools.
We find that the top four exclusive providers causing the highest proposer loss before 2024 are integrated searchers engaged in CEX-DEX arbitrage. This observation aligns with previous studies on competitive advantages and the profits associated with integration between these searchers and their integrated builders [61, 38], highlighting our algorithm’s effectiveness in identifying exclusive order flows.
Compared to the exclusive order flows before 2024, a more concerning problem is the formation of more exclusive relationships between top builders and top providers. This also confirms our conjecture. For instance, Banana Gun exclusively sends order flows to Titan in approximately 33K blocks, accounting for 37.1% of the estimated proposer losses covered in our study. Additionally, two other top-5 pivotal providers, jaredfromsubway.eth and Maestro, also exclusively send order flows to Beaver, contributing to 6% of the losses. In contrast, the total losses caused by the integrated CEX-DEX arbitrage searchers remain relatively similar before and during 2024. Notably, searcher 0x57c1… and searcher 0xf8b7… stopped operating in 2023. These results effectively explain that compared to the previous known integration, the emergence of new integration accounts for the majority of the proposer losses in 2024.
Finding IV: Top pivotal providers Banana Gun, jaredfromsubway.eth, and Maestro integrate with the top two builders, Titan and Beaver. Their integration causes significant losses to proposers.
Competing providers.   In addition to noting that three of the top five pivotal providers have started integrating with Beaver or Titan, we also observe that providers engaging in the same MEV activities tend to avoid integrating with the same builder simultaneously (Banana Gun sent order flows exclusively on Beaver in January 2024, while its flows to Titan occurred after this period). A possible explanation is concern over trust: if a provider knows that a builder collaborates with its competitor, it may worry about the security of its order flow sent to that builder. The builder could potentially leak information about the provider’s order flow to the competitor.
6 Discussion
6.1 Implications of Integration
Beyond proposer loss, the integration carries additional implications, which we discuss below.
Inaccurate MEV analysis.   Our study reveals a potential pattern between providers and builders: when the order flows from a provider may determine the auction outcome, the provider exclusively shares the order flow with a specific builder. In return, the provider may receive a refund from the builder. However, this refund could be off-chain, making it difficult to measure the providers’ actual costs. This limitation could lead to inaccuracies in MEV analysis tools. For example, the profit of a top MEV searcher, jaredfromsubway.eth, is only 1.3% of its total revenue according to [56]; its actual profit may be higher due to potential refunds.
Suboptimal block-building.   In principle, a builder with access to more profitable order flows can achieve better block-building. However, integration between top providers and builders can prevent builders from maximizing potential MEV, resulting in suboptimal block-building in the winners’ blocks and ultimately reducing profits for providers, builders, and proposers. For example, Beaver’s integrated searcher may identify a good opportunity for CEX-DEX arbitrage, while Banana Gun may also handle a large volume of private order flow exclusively sent to Titan. These two order flows are not in conflict; however, due to the builders’ competition, only one can be included in the winning block within the same slot. The excluded flow may not be included in the next block because the MEV opportunity could be time-sensitive (e.g., CEX-DEX arbitrages).
Censorship.   A security concern of builder centralization is transaction censorship. Ethereum’s current censorship resistance relies heavily on the choices of builders, particularly top builders. For example, before Beaver stopped censoring in July 2024, over 50% of blocks were built by censoring builders [62]. Our findings suggest that builder centralization remains difficult to overcome due to integration, potentially exacerbating the censorship problem on Ethereum.
6.2 Proposed Mitigations and Their Limitations
We now discuss why the proposed mitigation fails to resolve the identified problems. This highlights that addressing the identified problems is a grand challenge that requires further, possibly interdisciplinary, study.
Removing trust barriers.   There is a lack of a trustworthy mechanism to facilitate fair exchanges between private order flow providers and builders. As a result, some providers require builders to have a minimal market share, and we also observe that competing providers typically do not integrate with the same builder.
One promising direction to reduce this trust dependency is the use of encrypted mempools [63], where transactions remain encrypted until they are irreversibly included in a block, preventing builders from having complete knowledge of user transactions. For example, this can be implemented using Trusted Execution Environments (TEEs) [64] (e.g., SUAVE [65]), though standard challenges such as side channels and covert channels [66] still apply.
Besides, requiring builders to achieve fair ordering [25, 26], is another viable solution to prevent transaction manipulation, though a first-come-first-served ordering policy may introduce other problems like off-chain latency wars [67].
Proposed PBS designs.   The Ethereum community has considered multiple alternative designs to mitigate the centralization and security problems in the out-of-protocol PBS. Enshrined-PBS (ePBS) [68] is proposed to decentralize PBS auctions by removing centralized relays from PBS. MEV burn, an add-on of ePBS, which helps smooth and redistribute MEV spikes, benefits all ETH holders by burning the winning bid value in the PBS auction [13]. More recent proposals, including Execution Ticket [22] and Execution Auction [23], aim to further mitigate the effects of MEV on validator centralization by removing execution block proposing rights from validators.
However, in both eBPS and MEV burn, blocks are still produced through auctions, and the PBS auction can continue to exist as a secondary market in both Execution Ticket and Execution Auction. This allows the re-sale of block proposing rights to builders who can build more profitable blocks [22]. None of these proposals considers the incentives of providers, so the integration problem in PBS auctions remains unresolved.
Future work.   Based on the discussion, we identify that an effective mitigation must satisfy at least the following two properties: First, it must provide a security guarantee to providers’ order flows from potential attacks. Second, it should consider providers’ incentives, such as offering a game-theoretical guarantee to ensure they no longer have the motivation to continue such integration. Identifying these two properties is a step forward, but fully addressing them requires significant further study for future work.
6.3 Limitations of Our Approach and Data
We compute the true values of bids following a standard approach as [41, 42, 43, 44]. This approach relies on identifying builder-controlled addresses, and missing addresses could cause inaccuracies. We cross-validate with multiple sources to ensure collected addresses are the most comprehensive to date (see Sec. 3.1). It may underestimate values by ignoring off-chain payments, or overestimate profits if builders refund providers off-chain. A builder with large off-chain payments could bid higher than the computed true value, appearing as an extremely large subsidy. However, we observe that there are no excessively large subsidies, and most builders bid close to our computed true values in our study of competitiveness and inequality, suggesting that off-chain payments are not significant. The potential refunds to providers do not affect the computation of proposers’ losses because this portion is not allocated to proposers, regardless of the volume of refunds. Due to page constraints, we discuss the potential implications of the limitations in Appx. F.
@@ 7 Related Work
Empirical study on Ethereum’s builder market.   Several previous works studied the centralization of the builder market shares. An early exploration by Yang et al. [58] analyzed the market share of relays and builders. Their findings indicated a centralization of Ethereum’s builder market, with Flashbots and Builder0x69 accounting for over 53% of the market share between September 15th, 2022, and November 30th, 2022. Similar studies on the builder market conducted by [43, 69] also confirmed the centralization in the builder market. Their observations were based on the market share and the distribution of total profits generated within the PBS ecosystem. Moreover, several online dashboards [29, 70] offer a dynamic view of the ecosystem’s trends by presenting real-time PBS analytics. However, these studies do not quantify the causes of centralization or the security implications of a centralized builder market.
@@
Private order flow.   Many studies also discuss the importance of private order flows, particularly focusing on private order flows from integrated searchers. Thiery [44] counted the total number of transactions, along with their values, from June 1 to July 15, 2023. He found that private order flows account for only 30% of the transactions yet represent 80% of the total value paid to builders. Our study highlights a similar result: private order flows account for more than 80% of the block value in most blocks. Gupta et al. [20] confirmed that builders are more likely to win the MEV-Boost auctions when their integrated searchers can provide high-value exclusive order flow by effectively exploiting CEX-DEX arbitrage. Heimbach et al. [38] observed that during their study, Beaver received a total of 1,941.1 ETH in transaction fees but paid proposers 6,620.94 ETH. This large gap suggests that Beaver might have received significant off-chain profits from its integrated searchers. Öz et al. [57] employed Linear Discriminant Analysis to identify the integration between builders and providers. Our study uncovers previously unknown integration of top providers and quantifies the resulting proposer losses.
@@
Strategic behaviors in MEV-Boost auctions.   The strategies used by different entities in the MEV-Boost auctions and their potential impacts have attracted significant interest from researchers. The strategic behaviors of proposers were first studied by [69, 71] focusing on the timing game where proposers strategically delay block proposals to maximize their profits. Öz et al. [72] further conducted an agent-based simulation to analyze how waiting games affect consensus stability, finding that a delay strategy can be profitable and not degrade consensus if sufficiently many validators adopt it. Wu et al. [73] proposed a game-theoretic model for MEV-Boost auctions, using simulations to examine various builders’ bidding strategies, including naive, adaptive, last-minute, and bluff bidding. Their results highlighted the importance of latency and exclusive order flow in the effectiveness of bidding strategies. Pai et al. [9] analyzed the latency advantage of searcher-builder integration in adjusting bidding values. The strategies discussed in these papers highlight the potential factors that can make MEV-Boost auctions uncompetitive or inefficient, while our paper quantifies these uncompetitive and inefficient auctions.
@@
Causes and implications of centralized builder market.   Previous works such as [20, 74] theoretically analyzed the potential centralizing effects of exclusive order flows on the builder market: builders with exclusive order flows are more likely to win MEV-Boost auctions, thus gaining further advantages in receiving order flows from non-integrated searchers. Capponi et al. identified this dynamic as leading to a “Proof-Of-MEV” paradigm [75], where the capacity to capture MEV becomes a critical factor for success in MEV-Boost auctions. The decentralization of Ethereum’s builder market not only benefits itself but also positively impacts the decentralization of validators. A recent study showed that heterogeneity among validators in building blocks may lead to centralization; however, a decentralized and competitive builder market can encourage validator decentralization [5]. We complement these theoretical analyses with empirical studies to analyze the causes and potential problems of the current centralized builder market.